Calculate the acceleration of gravity relative to earth's acceleration (i.e., 1 g) at the surface of Jupiter. The mass of Jupiter is 1.9 x 10²⁷ kg and the planet's radius is 7.15 x 10⁷ m. The universal gravitational constant = G = 6.67 x 10⁻¹¹ N*m²/kg².


1
Expert's answer
2021-03-29T09:01:57-0400
"g_J=\dfrac{GM_J}{R_J^2},""g_J=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot1.9\cdot10^{27}\ kg}{(7.15\cdot10^7\ m)^2},""g_J=24.78\ \dfrac{m}{s^2}\cdot\dfrac{1\ g}{9.8\ \dfrac{m}{s^2}}=2.53g."